OpenStudy (twistrose):

1. Edward’s recipe for fudge uses a ratio of 2 1/4 cups of nuts to 1/2 lb of fudge. (a) Use a complex fraction to write a rate comparing the amount of nuts to the amount of fudge. (b) Use the unit rate to determine how many cups of nuts are needed to make 1 lb of fudge. Show your work.

OpenStudy (anonymous):

I think its b

OpenStudy (twistrose):

i needhelp with both parts. its two different parts

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

amount of nuts to one pound of fudge is \(4\tfrac{1}{2}\) to \(\frac{1}{2}\)

OpenStudy (misty1212):

oops i meant \[4\tfrac{1}{4}:\frac{1}{2}\]

OpenStudy (misty1212):

as a complex fraction that is \[\huge \frac{4\tfrac{1}{4}}{\frac{1}{2}}\]

OpenStudy (misty1212):

dividing by one half is the same as multiplying by two, so to finish multiply \[2\times 4\tfrac{1}{4}\]

OpenStudy (twistrose):

ok thanks

OpenStudy (anonymous):

oops